OpenStudy (anonymous):

trying to find the lateral surface area of a cylinder. the formula is S=2πrh. (h) height is 4", (d) diameter is 3". (r) radius is half of the diameter. how big is the area? not sure how pi figures in here. do you times them all? thanks!

OpenStudy (anonymous):

Yes, you multiply them all.

OpenStudy (anonymous):

Also please don't repeat questions when you don't get answers right away.

OpenStudy (anonymous):

do you know the answer? show your work if you could please.

OpenStudy (anonymous):

sorry, so I take 2 x π then take that times 1.5 then times 4? I get 37.7 which doesn't seem right?

OpenStudy (anonymous):

\[2\pi(1.5)(4) = 8(1.5)\pi = 12\pi \approx 37.7\] Yeah, that looks about right to me.